find three consecutive even integers such that the product of the first and the third integers equals 4 less than 8 times the second integer

Answer by CubeyThePenguin(3113)

consecutive even integers: (x-2), x, (x+2)

(x-2)(x+2) = 8x - 4
x^2 - 4 = 8x - 4
x^2 - 8x = 0
x(x - 8) = 0

The integers could be {-2, 0, 2} or {6, 8, 10.}